# Bayesian Adaptive Tucker Decompositions for Tensor Factorization

### v1

@@ 4.2 Licorice chemometrics data
    Chemical data are often multiway by nature, making tensor factorization techniques popular in chemometrics (Tomasi and Bro, 2005). In this section we analyze a sensor-based dataset from the licorice industry (Skov and Bro, 2005), driven by the food industry’s increasing focus on process control and quality checking. Data from sensor-based analytical instruments (e.g. electronic nose) can be arranged in a three-way array as sample $\times$ time $\times$ sensor. The electronic nose is used to distinguish complex volatiles, which can reproduce the structure and principle of olfactory sense. The datasets consists of $6$ good licorice samples and $12$ bad licorice samples, half of which are artificially fabricated by drying for a long time good samples. Twelve sensors, all based on Metal Oxide Semiconductor technologies, were used to capture the volatile compound measurements every 0.5 seconds over a span of 120 seconds, resulting in a data tensor $\mathcal{Y}$ of dimensions $18\times 241\times 12$. The sensors are distributed in two chambers with different temperatures. The data are normalized over all cells to have mean zero and unit variance, allowing one to assume default values for hyper-parameters in the proposed multiway priors.
@@ 4.3 Fish biodiversity data
    We analyze data from a community ecology monitoring program of the North Sea demersal fish community. The International Council for the Exploration of the Sea (ICES) created a unique data set covering more than 30 years of multiple fish species abundance across different locations in the North Sea. Data have been collected using standard otter trawl as sampling gear. The North Sea marine ecosystem has already been significantly impacted by climate change and fisheries exploitation, motivating a comprehensive analysis of fish community dynamics to support efforts in preserving this fragile ecosystem.
@@
    To investigate the spatial and temporal dynamics of fish communities we look at the posterior marginal probabilities $\pi_{\mathbf{i}}$ in (8) composing the corresponding probability tensor $\mathcal{M}$. Analyzing these probabilities across different tensor slices can provide valuable insights into population behavior, potentially revealing interesting spatial or temporal patterns. Figure S3 in the Supplementary Materials shows the posterior mean of the frontal slice of $\mathcal{M}$, $M_{::i}$ for $i=1,\dots,n_{3}$, for two roundfish areas of the North Sea: the northernmost area and one located in the southeast. The latter highlights the differences between fish species in the southern part of the North Sea and those in the northern region, which is deeper and experiences less seasonal temperature variation. As a result, we observe a strong north-south component in the composition of the North Sea fish community, consistent with previous findings (Frelat et al., 2017). Looking at the bio-geographic information of species, we observe that Atlantic fishes, species widespread in the North Atlantic, have a small probability of occurring in the southern area close to the Netherlands’coasts, while they have a higher probability of occurring in the northeast area, as it borders the Atlantic Ocean. Conversely, Lusitanian fishes have a higher probability of occurring in the southern area. Lusitanian fishes, indeed, tend to be abundant from the Iberian peninsula to as far north as the British Isles and the central North Sea (Engelhard et al., 2011).
@@ Acknowledgements
    The authors thank David Dunson and Alberto Cassese for comments on early versions of the draft.
@@
    We start by observing that
    $\mathrm{pr}\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}\leq\epsilon\}$ $=E[\mathrm{pr}\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}\leq\epsilon\mid\theta^{(1)},\dots,\theta^{(K)}\}]$
    $=1-E[\mathrm{pr}\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}>\epsilon\mid\theta^{(1)},\dots,\theta^{(K)}\}]$
    $=1-\frac{E[E\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}\mid\theta^{(1)},\dots,\theta^{(K)}\}]}{\epsilon}.$
    Hence we will focus on computing $E[E\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}\mid\theta^{(1)},\dots,\theta^{(K)}\}]=E\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}\}$. Since $g_{\mathbf{r}}$ and $\{u^{(k)}_{i_{k}r_{k}}\}_{k=1}^{K}$ are independent for all indexes $\mathbf{r}$ and $E(u^{(k)}_{i_{k}r_{k}})=0$, the latter expected value is equal to
    $\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}E\Big{\{}\Big{(}g_{\mathbf{r}}\prod_{k=1}^{K}u^{(k)}_{i_{k}r_{k}}\Big{)}^{2}\Big{\}}=\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}E(g_{\mathbf{r}}^{2})\prod_{k=1}^{K}E\{(u^{(k)}_{i_{k}r_{k}})^{2}\},$
    where $E\{(u^{(k)}_{i_{k}r_{k}})^{2}\}=E[E(\{u^{(k)}_{i_{k}r_{k}}\}^{2}\mid\theta^{(k)})]=E(1/\theta^{(k)}_{rk})=1/\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{r_{k}}$ as shown in Legramanti et al. (2020). Under the assumption that the elements of the core tensor $g_{\mathbf{r}}$ have bounded variances, $E(g_{\mathbf{r}}^{2})$ will be bounded by $\zeta<\infty$. Therefore, the following bound holds
    $\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}E\Big{\{}\Big{(}g_{\mathbf{r}}\prod_{k=1}^{K}u^{(k)}_{i_{k}r_{k}}\Big{)}^{2}\Big{\}}<\zeta\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}\prod_{k=1}^{K}\frac{1}{\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{r_{k}}}.$
    By noticing that $\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{r_{k}}=\alpha^{(k)}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{R_{k}}$ (Legramanti et al., 2020) it follows that
    $E[E\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}\mid\theta^{(1)},\dots,\theta^{(K)}\}]<\zeta\prod_{k=1}^{K}\frac{1}{\alpha^{(k)}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{R_{k}}},$
    from which follows that $\mathrm{pr}\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}>\epsilon\}<\zeta\prod_{k=1}^{K}\frac{1}{\epsilon\alpha^{(k)}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{R_{k}}}$. ∎
@@
[…]
    $\mathrm{pr}(E_{n}\mid y)=\frac{\int_{E_{n}}\prod_{i=1}^{n}\frac{f(y_{i})}{f_{0}(y_{i})}d\Pi(\eta_{i})}{\int\prod_{i=1}^{n}\frac{f(y_{i})}{f_{0}(y_{i})}d\Pi(\eta_{i})}=\frac{N_{n}}{D_{n}}.$ (12)
@@
    where $H=\{l=\tilde{R}^{0n}+1,\dots,n\}$. If we decompose the probability in (12) in the sum of two complementary conditional probabilities, it follows from the law of total probability that
@@
    Thanks to Lemma 5.2 of Castillo and van der Vaart (2012) $\mathrm{pr}_{0}(A_{n}^{c})\leq\exp(-r_{n}^{2})$ for increasing $n$, from which follows that using (12) we obtain
@@
    Using $(1-x)^{1/x}>1/(2e)$ for $0.50<x<$ 0.5 it follows that $\mathrm{pr}(||\eta_{H}||\leq r_{n})\geq(1/2e)^{KA\tilde{R}^{0n}}$. Then, choosing $r_{n}=\tilde{R}^{0n}$ it follows by Lemma 1 that
    $E_{0}\{\mathrm{pr}(E_{n}\mid y_{n})\}\leq\exp[-\tilde{R}^{0n}\{CK-1-KA\log(2e)\}]+\exp(-\tilde{R}^{0n}).$ (13)
    Since $C>2Ae$ implies $C>1+KA\log(2e)$ for $A>1/2$ and $K>1$, (13) yields $E_{0}\{\mathrm{pr}(E_{n}\mid y_{1},\dots,y_{n})\}\to 0$.
@@
    Under the multiway CUSP prior in (4) with $\theta_{\infty}=\theta_{\infty}(n)>r/\sqrt{n}$, for $1<r<\sqrt{n}$
@@
    In this section, we provide practical details on the implementation of the adaptive Gibbs sampler outlined in Algorithm 2 and discuss posterior inference for the probit-AdaTuck model. To let the chain stabilize, we start the adaptation after a fixed number of iterations $\tilde{t}$. In our analysis, we set $\tilde{t}=400$. The adaptation is performed with probability $p(t)=\exp(c_{0}+c_{1}t)$ and following Bhattacharya and Dunson (2011) we set $(c_{0},c_{1})=(-1,-5\times 10^{-4})$. It is reasonable to initialize $(R_{1},\dots,R_{K})$ with sufficiently large values such that $R_{k}\leq n_{k}+1$. We found in practice that initializing $R_{k}$ at $\{\mathrm{max}(n_{1},\dots,n_{k})+n_{k}\}/3$ for $k=1,\dots,K$ yields a good trade-off in terms of conservative truncation levels and computational scalability. Lastly, we initialize $R^{*}_{k}$ to $R_{k}-1$.
    Posterior inference for the probit-AdaTuck model defined in (7) of the main paper proceeds through an adaptive Gibbs sampler which exploits a data-augmentation scheme for the latent tensor $\mathcal{Q}$ (Albert and Chib, 1993). The adaptation strategy is the same as discussed in Section 3 of the main paper, so posterior inference still follows Algorithm 2. However, what differs for the probit-AdaTuck model are the Gibbs sampler’s sampling steps. Algorithm 3 outlines these specific steps for the probit-AdaTuck model. For dealing with missing entries we follow the same strategy discussed in the main paper, i.e we run the sampler conditioned only on the observed data. Specifically, we sample $q_{\mathbf{i}}$ from a univariate truncated normal (step 1 in Algorithm 3) only for the corresponding observed $y_{\mathbf{i}}$, resulting in a computationally efficient procedure.
@@
    S.3 Additional figures for illustrations

### v2

@@ 4.2 Licorice chemometrics data
    Chemical data are often multiway by nature, making tensor factorization techniques popular in chemometrics (Tomasi and Bro, 2005). In this section, we analyze a sensor-based dataset from the licorice industry (Skov and Bro, 2005), driven by the food industry’s increasing focus on process control and quality checking. Data from sensor-based analytical instruments (e.g. electronic nose) can be arranged in a three-way array as sample $\times$ time $\times$ sensor. The electronic nose is used to distinguish complex volatiles, which can reproduce the structure and principle of olfactory sense. The dataset consists of $6$ good licorice samples and $12$ bad licorice samples, half of which are artificially fabricated by drying for a long time good samples. Twelve sensors, all based on Metal Oxide Semiconductor technologies, were used to capture the volatile compound measurements every 0.5 seconds over a span of 120 seconds, resulting in a data tensor $\mathcal{Y}$ of dimensions $18\times 241\times 12$. The sensors are distributed in two chambers with different temperatures. The data are normalized over all cells to have mean zero and unit variance, allowing one to assume default values for hyperparameters in the proposed multiway priors.
@@ 4.3 Fish biodiversity data
    We analyze data from a community ecology monitoring program of the North Sea demersal fish community. The International Council for the Exploration of the Sea (ICES) created a unique data set covering more than 30 years of multiple fish species abundance across different locations in the North Sea. Data have been collected using standard otter trawls as sampling gear. The North Sea marine ecosystem has already been significantly impacted by climate change and fisheries exploitation, motivating a comprehensive analysis of fish community dynamics to support efforts in preserving this fragile ecosystem.
@@
    To investigate the spatial and temporal dynamics of fish communities we look at the posterior marginal probabilities $\pi_{\mathbf{i}}$ in (8) composing the corresponding probability tensor $\mathcal{M}$. Analyzing these probabilities across different tensor slices can provide valuable insights into population behavior, potentially revealing interesting spatial or temporal patterns. Figure S4 in the Supplementary Materials shows the posterior mean of the frontal slice of $\mathcal{M}$, $M_{::i}$ for $i=1,\dots,n_{3}$, for two roundfish areas of the North Sea: the northernmost area and one located in the southeast. The latter highlights the differences between fish species in the southern part of the North Sea and those in the northern region, which is deeper and experiences less seasonal temperature variation. As a result, we observe a strong north-south component in the composition of the North Sea fish community, consistent with previous findings (Frelat et al., 2017). Looking at the biogeographic information of species, we observe that Atlantic fishes, species widespread in the North Atlantic, have a small probability of occurring in the southern area close to the Netherlands ’ coasts, while they have a higher probability of occurring in the northeast area, as it borders the Atlantic Ocean. Conversely, Lusitanian fishes have a higher probability of occurring in the southern area. Lusitanian fishes, indeed, tend to be abundant from the Iberian peninsula to as far north as the British Isles and the central North Sea (Engelhard et al., 2011).
@@ Acknowledgements
    The authors thank David Dunson, Alberto Cassese and Giovanni Toto for comments on early versions of the draft.
@@
    Let $M_{\textbf{i}}^{R_{1},\dots,R_{K}}=\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}g_{\mathbf{r}}\prod_{k=1}^{K}u^{(k)}_{i_{k}r_{k}}$ be the residuals for a generic index vector $\mathbf{i}$. By Markov’s inequality
    $\mathrm{pr}\{(M_{\textbf{i}}^{R_{1},\dots,R_{K}})^{2}\geq\epsilon\}\leq\frac{E\{(M_{\textbf{i}}^{R_{1},\dots,R_{K}})^{2}\}}{\epsilon},$
    and since $g_{\mathbf{r}}$ and $\{u^{(k)}_{i_{k}r_{k}}\}_{k=1}^{K}$ are independent for all indexes $\mathbf{r}$, the second moment of $M_{\textbf{i}}^{R_{1},\dots,R_{K}}$ is equal to
    $E\{(M_{\textbf{i}}^{R_{1},\dots,R_{K}})^{2}\}=\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}E(g_{\mathbf{r}}^{2})\prod_{k=1}^{K}E\{(u^{(k)}_{i_{k}r_{k}})^{2}\},$
    where $E\{(u^{(k)}_{i_{k}r_{k}})^{2}\}=E[E\{(u^{(k)}_{i_{k}r_{k}})^{2}\mid\theta^{(k)}\}]$$={E(\theta^{(k)}_{r_{k}})}=\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{r_{k}}$ as shown in Legramanti et al. (2020). Under the assumption that the elements of the core tensor $g_{\mathbf{r}}$ have bounded variances, $E(g_{\mathbf{r}}^{2})$ will be bounded by $\zeta<\infty$. Therefore, the following bound holds
    $E\{(M_{\textbf{i}}^{R_{1},\dots,R_{K}})^{2}\}<\zeta\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}\prod_{k=1}^{K}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{r_{k}}.$
    The above equation can be factorized with respect to the indexes of the sum, and by noticing that $\sum_{r_{k}=R_{k}+1}^{\infty}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{r_{k}}=\alpha^{(k)}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{R_{k}}$ for any $k=1,\dots,K$ we obtain
    $E\{(M_{\textbf{i}}^{R_{1},\dots,R_{K}})^{2}\}<\zeta\prod_{k=1}^{K}\alpha^{(k)}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{R_{k}},$
    which concludes the proof.
@@
    We denote with $f(\cdot)$ the density of $y_{i}$ under model (9), while $f_{0}(\cdot)$ denotes the density of $y_{i}$ under the true model (10) in the main paper. Let $\Pi(\cdot)$ be the prior measure on $\eta_{i}=\prod_{k=1}^{K}u_{i}^{(k)}$. Let $E_{n}$ be the target event $E_{n}=\{M_{n}>\epsilon_{n}\}.$ Consistent with Castillo and van der Vaart (2012), we write
    $\mathrm{pr}(E_{n}\mid y)=\frac{\int_{E_{n}}\prod_{i=1}^{n}\frac{f(y_{i})}{f_{0}(y_{i})}d\Pi(\eta_{i})}{\int\prod_{i=1}^{n}\frac{f(y_{i})}{f_{0}(y_{i})}d\Pi(\eta_{i})}=\frac{N_{n}}{D_{n}}.$ (S.1)
@@
    where $H=\{l=\tilde{R}^{0n}+1,\dots,n\}$. If we decompose the probability in (S.1) in the sum of two complementary conditional probabilities, it follows from the law of total probability that
@@
    Thanks to Lemma 5.2 of Castillo and van der Vaart (2012) $\mathrm{pr}_{0}(A_{n}^{c})\leq\exp(-r_{n}^{2})$ for increasing $n$, from which follows that using (S.1) we obtain
@@
    Using $(1-x)^{1/x}>1/(2e)$ for $0.50<x<$ 0.5 it follows that $\mathrm{pr}(||\eta_{H}||\leq r_{n})\geq(1/2e)^{KA\tilde{R}^{0n}}$. Then, choosing $r_{n}^{2}=\tilde{R}^{0n}$ it follows by Lemma 1 that
    $E_{0}\{\mathrm{pr}(E_{n}\mid y_{n})\}\leq\exp[-\tilde{R}^{0n}\{CK-1-KA\log(2e)\}]+\exp(-\tilde{R}^{0n}).$ (S.2)
    Since $C>2Ae$ implies $C>1+KA\log(2e)$ for $A>1/2$ and $K>1$, (S.2) yields $E_{0}\{\mathrm{pr}(E_{n}\mid y_{1},\dots,y_{n})\}\to 0$.
@@
    Under the multiway CUSP prior in (4) in the main paper with $\theta_{\infty}=\theta_{\infty}(n)<(r/\sqrt{n})^{1/K}$, for $1<r<\sqrt{n}$ ,
@@
    In this section, we provide practical details on the implementation of the adaptive Gibbs sampler outlined in Algorithm 2 in the main paper and discuss posterior inference for the probit-AdaTuck model. To let the chain stabilize, we start the adaptation after a fixed number of iterations $\tilde{t}$. In our analysis, we set $\tilde{t}=400$. The adaptation is performed with probability $p(t)=\exp(c_{0}+c_{1}t)$ and following Bhattacharya and Dunson (2011) we set $(c_{0},c_{1})=(-1,-5\times 10^{-4})$. It is reasonable to initialize $(R_{1},\dots,R_{K})$ with sufficiently large values such that $R_{k}\leq n_{k}+1$. We found in practice that initializing $R_{k}$ at $\{\mathrm{max}(n_{1},\dots,n_{k})+n_{k}\}/3$ for $k=1,\dots,K$ yields a good trade-off in terms of conservative truncation levels and computational scalability. Lastly, we initialize $R^{*}_{k}$ to $R_{k}-1$.
    Posterior inference for the probit-AdaTuck model defined in (7) of the main paper proceeds through an adaptive Gibbs sampler which exploits a data-augmentation scheme for the latent tensor $\mathcal{Q}$ (Albert and Chib, 1993). The adaptation strategy is the same as discussed in Section 3 of the main paper, so posterior inference still follows Algorithm 2 in the main paper. However, what differs for the probit-AdaTuck model are the Gibbs sampler’s sampling steps. Algorithm 3 outlines these specific steps for the probit-AdaTuck model. For dealing with missing entries we follow the same strategy discussed in the main paper, i.e we run the sampler conditioned only on the observed data. Specifically, we sample $q_{\mathbf{i}}$ from a univariate truncated normal (step 1 in Algorithm 3) only for the corresponding observed $y_{\mathbf{i}}$, resulting in a computationally efficient procedure.
@@
    S.3 Image completion
    In this section, we analyze image data to assess AdaTuck’s performance in image reconstruction. A Red-Green-Blue (RGB) image can be naturally encoded in a three-way tensor, with dimension pixel height $\times$ pixel width $\times$ 3, where each entry corresponds to the intensity of the red, green, or blue color channels. The dataset used in our analysis is a cat image, depicted in panel (a) of Figure S1, with dimensions $120\times 107\times 3$. To assess the performance of AdaTuck in image reconstruction we randomly impute to missing values the $50\%$ and the $70\%$ of the total entries of the data $\mathcal{Y}$.
    Figure S1: Original image (a); image reconstructed by AdaTuck with $50\%$ missing values (b) and image reconstructed by AdaTuck with $70\%$ missing values (c).
    We apply AdaTuck to the image data, using the same hyperparameters and number of iterations for the Gibbs sampler described in Section 4.1 of the main paper. We benchmark the results with the Bayesian tensor factorization method proposed in Khan and Ammad-ud din (2016) (tensorBF). We compare the two methods by computing the Monte Carlo MSE in predicting the missing values with the reconstructed tensor based on the latent structure. Table S1 summarizes for the two scenarios the median and the interquartile range of the MSE obtained by AdaTuck and TensorBF. In both scenarios, AdaTuck outperforms the competitor in predicting the missing entries of the tensor. Figure S1 shows the cat image reconstructed by AdaTuck with $50\%$ missing values (panel b) and $70\%$ missing values (panel c).
    Table S1: Performance of AdaTuck and TensorBf in predicting the missing entries for the image data. The shorthand ‘NA(%)’ represents the percentage of missing entries with respect to the total tensor entries.
    AdaTuck TensorBF
    NA (%) MSE MSE MSE MSE
    (median) (IQR) (median) (IQR)
    50 0.012 0.003 0.014 0.002
    70 0.013 0.002 0.016 0.004
    S.4 Additional results for illustrations
    Table S2: Median and IQR of the latent multi-ranks estimated through the eigenvalue ratio (ER) and the Growth Ratio (GR) for the 20 different simulations in the two scenarios for $30\%$ and $50\%$ of total entries missing.
    ER GR
    $(R_{1}^{0},R_{2}^{0},R_{3}^{0})$ NA (%) $R_{1}$ $R_{2}$ $R_{3}$ $R_{1}$ $R_{2}$ $R_{3}$
    median(IQR) median(IQR)
    (5,5,5) 30 1(1) 1(1) 1(0.25) 1.5(2) 1(1) 1(1)
    50 1(0) 1(0) 1(1) 1(0.25) 1(0) 1(0.25)
    (10,7,3) 30 1(1) 1(1) 1(0) 1(1) 1.5(1) 1(0)
    50 1(1) 1(1) 1(0.25) 1(1) 1(1) 1(0)